# A Latent Variable Model with Change-Points and Its Application to Time Pressure Effects in Educational Assessment

### v2

@@ Abstract
Change-points,
Latent Factor Model,
Item Response Theory,
Educational Assessment,
keywords:
[…]
, , and
@@ 1 Introduction
Educational assessments play an important role for measuring student knowledge, skills, and growth in various academic domains. These tests are also used in educational decision-making, from informing instructional practices to high-stakes determinations such as college admissions. Given their importance, ensuring the validity and fairness of test score interpretations is paramount (see Chapter 3, American Educational Research Association et al., 2014, ). However, these measurements can be compromised when test takers exhibit changes in response behavior during the test due to, for example, time-pressure, fatigue or disengagement (Boughton and Yamamoto, 2007; Goegebeur et al., 2008; Wise and DeMars, 2006; Wang and Xu, 2015).
@@
In standardized educational testing, respondents are given a set of questions, referred to as items, constructed to measure the respondent’s ability levels in some well-defined domains. Multiple-choice items are typically used, with one correct response category per item. In the educational testing literature, such items are often referred to as binary items since the responses are scored as either correct or incorrect. A common assumption when analysing such testing data is that every respondent is giving every item full attention (Bolt, Cohen and Wollack, 2002; Schnipke and Scrams, 1997; Wise and Kong, 2005). In that way, the item responses can be assumed to only reflect the ability level of the respondent (plus some measurement error), given that the items are well-constructed. The probability of a correct answer therefore monotonically increases as the latent trait level increases (Wang and Xu, 2015). At the same time, standardized tests typically impose a time limit, terminating the test when the allotted time expires regardless of completion status. Due to time pressure, the imposed time limit can lead to a behavioral change, where the respondent responds without processing the item content (Wise and Kong, 2005). It is commonly known as test speededness in the educational testing literature. This change can be seen as a contamination of the signal of a latent variable: The respondent’s latent ability is the primary factor affecting the item responses up until the change, but thereafter, the responses start getting affected by time pressure effects as well. Post-change responses do therefore not accurately reflect the respondent’s true ability, which could lead to a potential misinterpretation of their performance.
@@
Item response theory (IRT) models have been extensively used in the development, assessment, and scoring of educational tests (Chen et al., 2021; Embretson and Reise, 2013). They provide a probabilistic framework that relates observed item responses to latent traits. IRT is the engine behind several educational testing innovations, such as (i) computerized adaptive testing (Van der Linden et al., 2000), which dynamically adjusts item difficulty to match a test-taker’s ability in real time, (ii) multistage testing (Yan, Von Davier and Lewis, 2016), which adjusts difficulty at predefined stages based on performance, (iii) automated test assembly (Linden et al., 2005), which uses algorithms to create assessments that meet specific statistical and content requirements, and (iv) personalized learning systems (Chen, Lee and Chen, 2005) which customize instruction, content, and pacing to individual student needs and abilities. At the core of all these frameworks is an IRT model.
@@
Traditional IRT models assume that all respondents give full attention to all items throughout the test, disregarding potential time-pressure effects or changes in problem-solving strategies (Van der Linden, 2007). In practice, test-taker behavior often deviates from this idealized assumption, particularly in timed testing scenarios. As test takers progress through an examination, factors such as fatigue, time constraints, or disengagement can lead to shifts in response patterns (Wise & DeMars, 2006). One well-documented phenomenon is “rapid guessing”, where respondents switch from careful problem-solving to quick, often random, responding as time runs out (Schnipke and Scrams, 1997). This behavior introduces construct-irrelevant variance into test scores, potentially biasing ability estimates and threatening the validity of inferences drawn from these assessments (Bolt, Cohen and Wollack, 2002).
@@
The detection and modeling of such behavioral changes during testing have received growing attention in psychometric research. Early approaches to identifying aberrant response patterns, such as person-fit statistics (Meijer and Sijtsma, 2001), provided global measures of response consistency but lacked the ability to pinpoint where changes in behavior occur. Other work has explored IRT models that allow for shifts in item parameters at a fixed point in the test, common to all test-takers (Bolt, Cohen and Wollack, 2002; Yamamoto and Everson, 1997). However, these models do not account for individual differences in the onset of behavioral changes, which may vary due to factors such as cognitive fatigue, time management skills, or the interaction between ability and item difficulty.
@@
To address these limitations, we turn to the literature on change-point detection. Change-point methods have been widely applied to detect abrupt shifts in data generating processes across various fields. These include, but are not limited to, bioinformatics (Olshen et al., 2004; Futschik et al., 2014), climatology (Reeves et al., 2007), distributed sensor networks (Yang, Eckley and Fearnhead, 2024), medicine (Bosc et al., 2003; Staudacher et al., 2005), entertainment (Rybach et al., 2009; Radke et al., 2005) and finance (Kim, Morley and Nelson, 2005; Andreou and Ghysels, 2002). We choose not to distinguish between online and offline settings, but instead point out that most of the developed change-point methods focus on either detecting a single change (e.g., Gombay and Horvath, 1990; Hinkley, 1970, 1971; Hawkins, 1977; James, James and Siegmund, 1987; Sen and Srivastava, 1975; Worsley, 1979), or multiple changes in a single data stream (e.g., Romano et al., 2022; Killick, Fearnhead and Eckley, 2012; Fryzlewicz, 2014; Fearnhead and Rigaill, 2019; Yao, 1988; Zou, Wang and Li, 2020; Niu, Hao and Zhang, 2016; Maidstone et al., 2017; Fryzlewicz, 2018, among many others). Of particular relevance to our work is the literature on change-point detection for multiple data streams (Chan, 2017; Chen and Zhang, 2015; Mei, 2010; Xie and Siegmund, 2013; Fellouris and Sokolov, 2016), where in our context, each respondent’s sequence of item responses represents a separate data stream. However, most literature on multi-stream change-point detection focuses on detection problems with a shared change-point among the data streams. Exceptions include Chen and Li (2023), Chen, Lee and Li (2022), and Lu, Chen and Li (2022). It is noteworthy that in our setting, we require a method capable of detecting changes for each data stream at potentially different locations. Additionally, existing change-point methods do not take the complex interdependencies in IRT models into account. In IRT, item parameters and person abilities are simultaneously estimated, creating a multidimensional problem where changes in one component can affect all others. This interdependence requires the development of new methodologies that can simultaneously estimate item parameters, person abilities, and individual-specific change-points within the IRT framework.
@@
The present research addresses this challenge by proposing a new class of IRT models that allow for individual-level change-points in response behavior. Specifically, we extend the standard two-parameter logistic (2-PL) IRT model (Birnbaum, 1968) to allow item-specific parameters that carry an application-related interpretation to change after an item that varies by person. The location of the change-point is treated as a latent random variable, enabling us to estimate its distribution in the test-taker population. We develop an empirical Bayes estimation approach based on the marginal maximum likelihood function to simultaneously estimate item parameters, person abilities, and change-point locations.
@@
Our work contributes to an emerging line of psychometric research on learning and behavioral dynamics in assessment. Our contribution adds to a series of recent studies that have explored extensions of IRT models to capture within-person changes from aberrant behavior due to e.g., time presure, fatigue, or cheating. Wang and Xu (2015) proposes a mixture hierarchical model that uses both item responses and response time to model effects of changes in response style, Wang et al. (2018) adjusts the model in Wang and Xu (2015) to include item-level response-style effect parameters, and Wang, Xu and Shang (2018) combines a mixture modeling approach with a residual-based outlier detection method to distinguish normal behavior from aberrant behavior. Similar to our approach, Shao, Li and Cheng (2016) formulates the person-level change as a change-point detection problem, but assumes all of the other model parameters to be known. Our proposal contributes to this stream of research by allowing for qualitative shifts in response behavior at a respondent-unique position while simultaneously estimating all the model parameters. This moves beyond static conceptions of ability and response behavior to provide a richer understanding of test-taker performance.
@@ 2.2 Baseline IRT Measurement Model
is a reparameterization of the 2-Parameter Logistic (2-PL) model (Birnbaum, 1968). If the slope parameters $a_{j}$, $j=1,\ldots,J$, are constrained to be constant and equal to 1, the resulting model is a reparameterization of the popular Rasch model (Rasch, 1960),
@@ 3.1 Marginal Likelihood
The inference of the proposed model is based on the marginal likelihood function (Bock and Aitkin, 1981), where the item parameters $\boldsymbol{\eta}_{1}$ and the distribution of the person parameters $p(\theta,\tau;\boldsymbol{\eta}_{2})$ are estimated simultaneously. Assuming that the responses $Y_{ij}$ are conditionally independent across items $j$ for each individual $i$ given the latent trait $\theta_{i}$ and change-point $\tau_{i}$ (known as local independence), the marginal maximum likelihood estimator is given by
@@
The term $p(\tau_{i}\mid\boldsymbol{\eta}_{2})$ denotes the probability mass function of the change-point $\tau_{i}$, and $p(\theta\mid\boldsymbol{\eta}_{2})$ represents the density function of the latent trait $\theta_{i}$. The vector $\boldsymbol{\eta}_{1}$ includes the item-specific parameters while $\boldsymbol{\eta}_{2}$ comprises the structural parameters governing the distributions of $\theta$ and $\tau$. As explained in Section 2.4, a parametric form of both $\theta$ and $\tau$ is assumed. The estimator based on (3.1) can be viewed as an empirical Bayes estimator (Chen et al., 2021).
@@
Given the complexity of the marginal likelihood function, which involves both integration over the latent trait $\theta$ and summation over the change-point $\tau$, a closed-form analytical solution does not exist. Therefore, the maximization of the marginal likelihood function must be carried out using iterative numerical optimization methods. Various approaches can be employed for this purpose, including the Expectation-Maximization (EM) algorithm (Bock and Aitkin, 1981; Dempster, Laird and Rubin, 1977), which is widely used in latent variable models. In our implementation, we utilize a quasi-Newton optimization method.
@@ 3.3 Model Selection
    For each model $M_{k}$, compute an information criterion $IC(M_{k})$. While various criteria exist, the Bayesian Information Criterion (BIC; Schwarz, 1978) is often preferred due to its consistency properties in model selection.
@@ 4.3 Extension to Polytomous Items
The model can be generalized to accommodate polytomous items using, for instance, the Graded Response Model (GRM, Samejima, 1969). For an item $j$ with $K_{j}$ ordered categories, the GRM can be reparameterized to express the cumulative probability of responding in category $k$ or higher as:
@@ 6.1 Dataset 1
(a) Estimated marginal distribution of $\tau$ for the first empirical dataset.
(b) Estimated marginal distribution of $\tau$ for the second empirical dataset.
@@ 7 Discussion
A strength of our modeling approach lies in its explicit incorporation of a temporal structure through the discrete-time hazard model, in contrast to approaches such as in Wang and Xu (2015) that only models the presence or absence of changes through a Bernoulli distribution. While such models can identify whether a change occurred, our framework provides richer insights by modeling when these changes are likely to occur. This temporal modeling is particularly valuable in educational testing contexts, where the timing of behavioral changes may carry important diagnostic information - for example, distinguishing between early changes that might reflect students’ initial adjustment to the testing environment and later changes that could indicate speededness effects. The model’s parameters ($\alpha$ and $\beta$) allow researchers and practitioners to better understand the dynamics of test-taking behavior. This approach enables not only the detection of behavioral changes but also insights into their temporal distribution, which can be valuable for test development and administration.

### v3

# inkscapelatex=false

@@ Abstract
Keywords: Change-points; Latent Factor Model; Item Response Theory; Educational Assessment.
@@ 1 Introduction
Educational assessments play an important role for measuring student knowledge, skills, and growth in various academic domains. These tests are also used in educational decision-making, from informing instructional practices to high-stakes determinations such as college admissions. Given their importance, ensuring the validity and fairness of test score interpretations is paramount (see Chapter 3, American Educational Research Association et al., 2014, ). However, these measurements can be compromised when test takers exhibit changes in response behavior during the test due to, for example, time-pressure, fatigue or disengagement (Boughton and Yamamoto, , 2007; Goegebeur et al., , 2008; Wise and DeMars, , 2006; Wang and Xu, , 2015).
@@
In standardized educational testing, respondents are given a set of questions, referred to as items, constructed to measure the respondent’s ability levels in some well-defined domains. Multiple-choice items are typically used, with one correct response category per item. In the educational testing literature, such items are often referred to as binary items since the responses are scored as either correct or incorrect. A common assumption when analysing such testing data is that every respondent is giving every item full attention (Bolt et al., , 2002; Schnipke and Scrams, , 1997; Wise and Kong, , 2005). In that way, the item responses can be assumed to only reflect the ability level of the respondent (plus some measurement error), given that the items are well-constructed. The probability of a correct answer therefore monotonically increases as the latent trait level increases (Wang and Xu, , 2015). At the same time, standardized tests typically impose a time limit, terminating the test when the allotted time expires regardless of completion status. Due to time pressure, the imposed time limit can lead to a behavioral change, where the respondent responds without processing the item content (Wise and Kong, , 2005). It is commonly known as test speededness in the educational testing literature. This change can be seen as a contamination of the signal of a latent variable: The respondent’s latent ability is the primary factor affecting the item responses up until the change, but thereafter, the responses start getting affected by time pressure effects as well. Post-change responses do therefore not accurately reflect the respondent’s true ability, which could lead to a potential misinterpretation of their performance.
@@
Item response theory (IRT) models have been extensively used in the development, assessment, and scoring of educational tests (Chen et al., , 2021; Embretson and Reise, , 2013). They provide a probabilistic framework that relates observed item responses to latent traits. IRT is the engine behind several educational testing innovations, such as (i) computerized adaptive testing (Van der Linden et al., , 2000), which dynamically adjusts item difficulty to match a test-taker’s ability in real time, (ii) multistage testing (Yan et al., , 2016), which adjusts difficulty at predefined stages based on performance, (iii) automated test assembly (Linden et al., , 2005), which uses algorithms to create assessments that meet specific statistical and content requirements, and (iv) personalized learning systems (Chen et al., , 2005) which customize instruction, content, and pacing to individual student needs and abilities. At the core of all these frameworks is an IRT model.
@@
Traditional IRT models assume that all respondents give full attention to all items throughout the test, disregarding potential time-pressure effects or changes in problem-solving strategies (Van der Linden, , 2007). In practice, test-taker behavior often deviates from this idealized assumption, particularly in timed testing scenarios. As test takers progress through an examination, factors such as fatigue, time constraints, or disengagement can lead to shifts in response patterns (Wise & DeMars, 2006). One well-documented phenomenon is “rapid guessing”, where respondents switch from careful problem-solving to quick, often random, responding as time runs out (Schnipke and Scrams, , 1997). This behavior introduces construct-irrelevant variance into test scores, potentially biasing ability estimates and threatening the validity of inferences drawn from these assessments (Bolt et al., , 2002).
@@
The detection and modeling of such behavioral changes during testing have received growing attention in psychometric research. Early approaches to identifying aberrant response patterns, such as person-fit statistics (Meijer and Sijtsma, , 2001), provided global measures of response consistency but lacked the ability to pinpoint where changes in behavior occur. Other work has explored IRT models that allow for shifts in item parameters at a fixed point in the test, common to all test-takers (Bolt et al., , 2002; Yamamoto and Everson, , 1997). However, these models do not account for individual differences in the onset of behavioral changes, which may vary due to factors such as cognitive fatigue, time management skills, or the interaction between ability and item difficulty.
@@
To address these limitations, we turn to the literature on change-point detection. Change-point methods have been widely applied to detect abrupt shifts in data generating processes across various fields. These include, but are not limited to, bioinformatics (Olshen et al., , 2004; Futschik et al., , 2014), climatology (Reeves et al., , 2007), distributed sensor networks (Yang et al., , 2024), medicine (Bosc et al., , 2003; Staudacher et al., , 2005), entertainment (Rybach et al., , 2009; Radke et al., , 2005) and finance (Kim et al., , 2005; Andreou and Ghysels, , 2002). We choose not to distinguish between online and offline settings, but instead point out that most of the developed change-point methods focus on either detecting a single change (e.g., Gombay and Horvath, , 1990; Hinkley, , 1970, 1971; Hawkins, , 1977; James et al., , 1987; Sen and Srivastava, , 1975; Worsley, , 1979), or multiple changes in a single data stream (e.g., Romano et al., , 2022; Killick et al., , 2012; Fryzlewicz, , 2014; Fearnhead and Rigaill, , 2019; Yao, , 1988; Zou et al., , 2020; Niu et al., , 2016; Maidstone et al., , 2017; Fryzlewicz, , 2018, among many others). Of particular relevance to our work is the literature on change-point detection for multiple data streams (Chan, , 2017; Chen and Zhang, , 2015; Mei, , 2010; Xie and Siegmund, , 2013; Fellouris and Sokolov, , 2016), where in our context, each respondent’s sequence of item responses represents a separate data stream. However, most literature on multi-stream change-point detection focuses on detection problems with a shared change-point among the data streams. Exceptions include Chen and Li, (2023), Chen et al., (2022), and Lu et al., (2022). It is noteworthy that in our setting, we require a method capable of detecting changes for each data stream at potentially different locations. Additionally, existing change-point methods do not take the complex interdependencies in IRT models into account. In IRT, item parameters and person abilities are simultaneously estimated, creating a multidimensional problem where changes in one component can affect all others. This interdependence requires the development of new methodologies that can simultaneously estimate item parameters, person abilities, and individual-specific change-points within the IRT framework.
@@
The present research addresses this challenge by proposing a new class of IRT models that allow for individual-level change-points in response behavior. Specifically, we extend the standard two-parameter logistic (2-PL) IRT model (Birnbaum, , 1968) to allow item-specific parameters that carry an application-related interpretation to change after an item that varies by person. The location of the change-point is treated as a latent random variable, enabling us to estimate its distribution in the test-taker population. We develop an empirical Bayes estimation approach based on the marginal maximum likelihood function to simultaneously estimate item parameters, person abilities, and change-point locations.
@@
Our work contributes to an emerging line of psychometric research on learning and behavioral dynamics in assessment. Our contribution adds to a series of recent studies that have explored extensions of IRT models to capture within-person changes from aberrant behavior due to e.g., time presure, fatigue, or cheating. Wang and Xu, (2015) proposes a mixture hierarchical model that uses both item responses and response time to model effects of changes in response style, Wang et al., 2018b adjusts the model in Wang and Xu, (2015) to include item-level response-style effect parameters, and Wang et al., 2018a combines a mixture modeling approach with a residual-based outlier detection method to distinguish normal behavior from aberrant behavior. Similar to our approach, Shao et al., (2016) formulates the person-level change as a change-point detection problem, but assumes all of the other model parameters to be known. Our proposal contributes to this stream of research by allowing for qualitative shifts in response behavior at a respondent-unique position while simultaneously estimating all the model parameters. This moves beyond static conceptions of ability and response behavior to provide a richer understanding of test-taker performance.
@@ 2.2 Baseline IRT Measurement Model
is a reparameterization of the 2-Parameter Logistic (2-PL) model (Birnbaum, , 1968). If the slope parameters $a_{j}$, $j=1,\ldots,J$, are constrained to be constant and equal to 1, the resulting model is a reparameterization of the popular Rasch model (Rasch, , 1960),
@@ 3.1 Marginal Likelihood
The inference of the proposed model is based on the marginal likelihood function (Bock and Aitkin, , 1981), where the item parameters $\boldsymbol{\eta}_{1}$ and the distribution of the person parameters $p(\theta,\tau;\boldsymbol{\eta}_{2})$ are estimated simultaneously. Assuming that the responses $Y_{ij}$ are conditionally independent across items $j$ for each individual $i$ given the latent trait $\theta_{i}$ and change-point $\tau_{i}$ (known as local independence), the marginal maximum likelihood estimator is given by
@@
The term $p(\tau_{i}\mid\boldsymbol{\eta}_{2})$ denotes the probability mass function of the change-point $\tau_{i}$, and $p(\theta\mid\boldsymbol{\eta}_{2})$ represents the density function of the latent trait $\theta_{i}$. The vector $\boldsymbol{\eta}_{1}$ includes the item-specific parameters while $\boldsymbol{\eta}_{2}$ comprises the structural parameters governing the distributions of $\theta$ and $\tau$. As explained in Section 2.4, a parametric form of both $\theta$ and $\tau$ is assumed. The estimator based on (3.1) can be viewed as an empirical Bayes estimator (Chen et al., , 2021).
@@
Given the complexity of the marginal likelihood function, which involves both integration over the latent trait $\theta$ and summation over the change-point $\tau$, a closed-form analytical solution does not exist. Therefore, the maximization of the marginal likelihood function must be carried out using iterative numerical optimization methods. Various approaches can be employed for this purpose, including the Expectation-Maximization (EM) algorithm (Bock and Aitkin, , 1981; Dempster et al., , 1977), which is widely used in latent variable models. In our implementation, we utilize a quasi-Newton optimization method.
@@ 3.3 Model Selection
    For each model $M_{k}$, compute an information criterion $IC(M_{k})$. While various criteria exist, the Bayesian Information Criterion (BIC; Schwarz, , 1978) is often preferred due to its consistency properties in model selection.
@@ 4.3 Extension to Polytomous Items
The model can be generalized to accommodate polytomous items using, for instance, the Graded Response Model (GRM, Samejima, , 1969). For an item $j$ with $K_{j}$ ordered categories, the GRM can be reparameterized to express the cumulative probability of responding in category $k$ or higher as:
@@ 6.1 Dataset 1
((a)) Estimated marginal distribution of $\tau$ for the first empirical dataset.
((b)) Estimated marginal distribution of $\tau$ for the second empirical dataset.
@@ 7 Discussion
A strength of our modeling approach lies in its explicit incorporation of a temporal structure through the discrete-time hazard model, in contrast to approaches such as in Wang and Xu, (2015) that only models the presence or absence of changes through a Bernoulli distribution. While such models can identify whether a change occurred, our framework provides richer insights by modeling when these changes are likely to occur. This temporal modeling is particularly valuable in educational testing contexts, where the timing of behavioral changes may carry important diagnostic information - for example, distinguishing between early changes that might reflect students’ initial adjustment to the testing environment and later changes that could indicate speededness effects. The model’s parameters ($\alpha$ and $\beta$) allow researchers and practitioners to better understand the dynamics of test-taking behavior. This approach enables not only the detection of behavioral changes but also insights into their temporal distribution, which can be valuable for test development and administration.